Find the standard form of the equation of the parabola with the given characteristic and vertex at the origin.
a. y^2=-x
b. y^2=-4x
c. x^2=-4y
d. x^2=-y
e. x^2=y

Answer by lwsshak3(11628)

Find the standard form of the equation of the parabola with the given characteristic and vertex at the origin.
a. y^2=-x
b. y^2=-4x
c. x^2=-4y
d. x^2=-y
e. x^2=y
**
standard form of a parabola: y=(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
If the coordinates of the vertex are at the origin, (0,0), as given, they do not appear in the equation
which becomes y=x^2.
Therefore, e, is the correct answer